OpenStudy (anonymous):

okay I meant to put square root of 4n to the third power? Would it be 2 square root n to the 3rd power?

OpenStudy (anonymous):

\[4^3=2^6=?\]

OpenStudy (anonymous):

\[2^3=8,2^4=16,2^5=32,2^6=\ldots?\]

OpenStudy (radar):

4nX4nX4n or \[4n ^{3}\]

OpenStudy (anonymous):

Oh, lol, I thought "nto" was a typo. In which case what radar said.

OpenStudy (radar):

I am not sure, i listed 2 possibilities: 64n^3 or 4n^3

OpenStudy (radar):

@jreynolds, please clarify is it (4n)^3 or is it 4n^3 ??

OpenStudy (radar):

Or is it something completely different lol.

OpenStudy (anonymous):

okay I meant to put square root of 4n to the third power? Would it be 2 square root n to the 3rd power?

OpenStudy (radar):

\[\sqrt{4n ^{3}}\]or\[\sqrt{(4n)^{3}}\]????

OpenStudy (radar):

For the first it would be \[2n \sqrt{n}\]for the second it would be\[2n \sqrt{2n}\]